1. Use any appropriate method to find each of the following integrals psin’ x-cos? 1.1 dx sin x+cos x [4]
Answer #1

sin^2 x-cos^2x is same as  a^2 -b^2=(a+b)(a-b)

Apply this in integral term and simply it

Now, integral become,

\int (sinx-cosx)dx

\int sinxdx= -cosx+c

\int cosxdx= sinx+c

Answer : -cosx- sinx+c

sin²x - cos²x dic sinxt loss s(sinsly co .dx [2²-6² = (atbla-b (singly closx) (sinx Cosy) (sire tlos x) = Slain x - Los x) du